# On Geometry and Topology of Hessian Manifolds

### v3

@@ 3.6 Riemannian Geometric Aspect
Let $((M,\nabla),h)$ be a compact Hessian manifold. Then for any Riemannian metric $g$ on $M$, the Ricci tensor of $(M,g)$ is not everywhere positive.
@@
If the Ricci tensor of $(M,g)$ is positive, then by Bonnet–Myers theorem, the fundamental group of $M$ is finite, which contradicts Corollary2.7. ∎
@@
3.7 Hesse’s Uncertainty Principle
Throughout this section, we use Einstein’s summation convention.
Lemma 3.61.
$\varkappa(Y)=\frac{1}{2}\sum_{i=1}^{n}A(e_{i},e_{i},Y)$
holds in orthonormal frame fields $e_{1},\dots,e_{n}$ of $(M,h)$.
Proof.
$2\varkappa_{k}Y^{k}=h^{ij}\partial_{k}h_{ij}Y^{k}=h^{ij}A_{ijk}Y^{k}$
as required. ∎
Lemma 3.62.
$\operatorname{div}_{h}(\eta^{\sharp})=-n+\varkappa(\eta^{\sharp})$
holds, where $\sharp$ is the musical isomorphism induce by $h$.
Proof.
Denote by $A$ the Amari–Chentsov tensor of $((M,\nabla),h)$ and denote by
$\bar{\nabla}:=\frac{1}{2}(\nabla+\nabla^{*})$
the Levi-Civita connection of $(M,h)$. Then, by definition, we have
$(\bar{\nabla}_{X}\eta)(Y)=h(X,Y)-\frac{1}{2}A(X,Y,\eta^{\sharp}).$
Therefore, by Lemma3.61, we conclude that
$\operatorname{div}_{h}(\eta^{\sharp})=-\star d\star\eta=-n+\varkappa(\eta^{\sharp})$
where $\star$ is the Hodge star operator on $(M,h)$. ∎
Corollary 3.63.
$\operatorname{Vol}_{h}(M)=\frac{1}{n}\int_{M}\varkappa\wedge\star\eta$
holds, where $\star$ is the Hodge star operator on $(M,h)$.
Proof.
Apply divergence theorem to the formula in Lemma3.62. ∎
Corollary 3.64.
$\operatorname{Vol}_{h}(M)^{2}\leq\frac{1}{n^{2}}\int_{M}|\varkappa|^{2}\operatorname{vol}_{h}\int_{M}|\eta|^{2}\operatorname{vol}_{h}$
holds on $M$.
Proof.
Apply Cauchy-Schwarz inequality to the integral formula in Corollary3.63. ∎
Remark 3.65.
The integral inequality stated in Corollary3.64 is an analog of Heisenberg uncertainty principle in quantum mechanics: For a given dimension and volume, a $1$-form potential small in $L^{2}$-norm must be compensated by a first Koszul form large in $L^{2}$-norm, and vice-versa. This phenomenon indicates a highly non-trivial relation between a $1$-form potential and the first Koszul form of a Hessian manifold. We name this integral inequality as Hesse’s uncertainty principle, in honor of the great German mathematician Ludwig Otto Hesse.
@@ 3.8 The Cheng-Yau Metric
Differentiate both sides of equation(1) again, we obtain that
@@
holds for all $\textup{{x}}\in\Omega$. Substitute equation(2) back into equation(1), we finally arrive at $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv 1$ as required. ∎
@@
This is nothing but a coordinate-dependent version of the formula in Lemma3.31. ∎
@@ 4.1 Complete Hessian surfaces
4.2 Compact Radiant Examples of Koszul Type
@@
for every $z\in\mathbb{C}$, we obtain that
$Q(\Phi(\gamma\cdot\varphi^{n}(\tau),\rho+n))=Q(\Phi(\tau,\rho))e^{2n}$
and hence
@@
The examples presented in this section confirm the previous Corollary2.44 in section2.
Remark 4.8.
The example above is consistent with the results established earlier. As predicted by Theorem2.41, for a closed orientable surface $\Sigma_{g}$ of genus $g>1$, the mapping torus $M_{\varphi}$ of a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ is indeed a fiber bundle over $\mathbb{S}^{1}$. The fundamental group of $M_{\varphi}$ is a semidirect product of $\mathbb{Z}$ and $\pi_{1}(\Sigma_{g})$, which is indeed infinite and torsion free, as stated in Proposition2.7 and Proposition2.19.
Remark 4.9.
It is obvious that the product of two Hessian manifolds is naturally a Hessian manifold. However, the converse statement is far from being correct, as in the example above, the product of the circle $\mathbb{S}^{1}$ and a closed orientable surface $\Sigma_{g}$ of genus $g>1$ admits an affine structure, but $\Sigma_{g}$ itself does not admit any affine structure if $g\neq 1$ by the proof of Chern’s conjecture in dimension two [20].
Therefore, in particular, the examples presented in this section are different to those constructed using Example1.7 and Theorem1.8.
4.3 Exotic Potentials for the Euclidean plane
Lemma 4.10.
Let $\Omega$ be a domain in $\mathbb{R}^{2}$, and let
Proof.
$F(x,y,z):=f(x/r,y/r,z/r)$
Proposition 4.11.
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
Proof.
Combine Corollary3.53 with Lemma4.10. ∎
Example 4.12.
$(2-d)H+x\frac{\partial H}{\partial x}+y\frac{\partial H}{\partial y}=0$
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a flat pseudo-Riemannian metric.
The following interesting example shows that nonlinear harmonic functions are Hessian potentials for the flat $1+1$ Minkowski plane.
Example 4.13.
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
Example 4.14.
$f(x,y):=\frac{x^{2}}{2y}+\frac{1}{4}\operatorname{log}(y)y.$
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
$f_{xx}^{2}-4f_{xx}f_{yy}+4f_{xy}^{2}\equiv 0$
and then apply Proposition4.11.
Before we close this section, we cite from [30] an example given by Dubrovin, to show that techniques from algebraic geometry, for example the Clebsch covariant, might be useful for finding homogeneous potentials.
Example 4.15.
of $f$ is a flat metric on $\mathbb{R}^{3}$.

### v4

@@ 4.1 Effects on Curvature Tensors
Let $((M,\nabla),h)$ be a complete Hessian manifold. Then for any Riemannian metric $g$ on $M$, then there exists no positive real number $k>0$ such that the Ricci tensor $\operatorname{Ric}^{g}$ of $(M,g)$ satisfies $\operatorname{Ric}^{g}\geq k$.
@@
If the Ricci tensor satisfies $\operatorname{Ric}^{g}\geq k$ for some positive real number $k>0$, then by Bonnet–Myers theorem, $M$ is compact with finite fundamental group, which contradicts Corollary 2.7. ∎
Corollary 4.7.
Let $((M,\nabla),g)$ be a compact Hessian manifold. Then for any Riemannian metric $h$ on $M$, the Riemannian manifold $(M,h)$ is not of positive sectional curvature.
Proof.
This is a direct consequence of Proposition 4.6. ∎
@@
Corollary 4.9 successfully shows that a Hessian metric on a compact manifold cannot be of negative sectional curvature. But, actually, using technique of harmonic mappings and Bochner formula, we can further prove that a compact Hessian manifold cannot support any metric of negative sectional curvature.
Theorem 4.10.
Let $((M,\nabla),g)$ be a compact Hessian manifold of dimension $n\geq 2$. Then for any Riemannian metric $h$ on $M$, the Riemannian manifold $(M,h)$ is not of negative sectional curvature.
Proof.
Assume the contrary that $h$ is a Riemannian metric on $M$ such that $(M,h)$ is of negative sectional curvature.
Suppose that $(M,g)$ is a flat Riemannian manifold. By Bieberbach theorem, there exists a finite covering map $\pi\colon\mathbb{T}^{n}\rightarrow M$. Therefore $(\mathbb{T}^{n},\pi^{*}h)$ is of negative sectional curvature. But then by Preissman’s theorem every non-trivial abelian subgroup of $\pi_{1}(\mathbb{T}^{n})=\mathbb{Z}^{n}$ is isomorphic to the inifinte cyclic group, which is absurd as $n\geq 2$.
From now on, we assume that $(M,g)$ is not a flat Riemannian manifold. Since $(M,h)$ is of non-positive sectional curvature, by Eells-Sampson theorem, there exists a harmonic map $f$ from $(M,g)$ to $(M,h)$ of mapping degree 1 homotopic to the identity map, and hence is in particular surjective.
Denote by $\nabla^{g}$ the Levi-Civita connection of $(M,g)$, and denote by $\varkappa\in\Omega^{1}(M)$ the first Koszul form of $((M,\nabla),g)$. Then by Proposition 2.43 we have $\nabla^{g}\varkappa=0$. Since $(M,h)$ is a compact Riemannian manifold of negative sectional curvature, we obtain that $f_{*}(\varkappa^{\sharp})\equiv 0$ by Theorem B.2.
By Proposition 2.44, we have that $\varkappa^{\sharp}$ is nowhere vanishing as $(M,g)$ is not flat. Therefore $\operatorname{dim}\operatorname{Ker}(df)\geq 1$, i.e. the differential $df$ is everywhere degenerated, and hence $f$ cannot be surjective by Sard’s Lemma, which contradicts the fact that $\operatorname{deg}(f)=1$.
This concludes the proof. ∎
The readers are invited to compare carefully Corollary 4.7 with Theorem 4.10.
@@ 4.2 The Cheng-Yau Metric
Differentiate both sides of equation( 1) again, we obtain that
@@
holds for all $\textup{{x}}\in\Omega$. Substitute equation( 2) back into equation( 1), we finally arrive at $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv 1$ as required. ∎
@@
This is nothing else but a coordinate-dependent and local version of the formula in Lemma 3.33. ∎
@@
Proposition 4.19 motivates the following more general definition.
Definition 4.20.
Let $((M,\nabla),h)$ be a Hessian manifold, and let $\varkappa\in\Omega^{1}(M)$ be its first Koszul form. The metric $h$ is said to be a Cheng-Yau metric if there exists $\eta\in\Omega^{1}(M)$ and a real constant $k>0$ such that $h=\nabla\eta$ and $\varkappa=k\eta$.
4.3 Koszul’s Uncertainty Principle
Throughout this section, we use Einstein’s summation convention.
Lemma 4.21.
$\varkappa(Y)=\frac{1}{2}\sum_{i=1}^{n}A(e_{i},e_{i},Y)$
holds in orthonormal frame fields $e_{1},\dots,e_{n}$ of $(M,h)$.
Proof.
$2\varkappa_{k}Y^{k}=h^{ij}\partial_{k}h_{ij}Y^{k}=h^{ij}A_{ijk}Y^{k}$
as required. ∎
Lemma 4.22.
$\operatorname{div}_{h}(\eta^{\sharp})=-\varkappa(\eta^{\sharp})+n$
holds, where $\sharp$ is the musical isomorphism induce by $h$.
Proof.
Denote by $A$ the Amari–Chentsov tensor of $((M,\nabla),h)$ and denote by
$\bar{\nabla}:=\frac{1}{2}(\nabla+\nabla^{*})$
the Levi-Civita connection of $(M,h)$. Then, by definition, we have
$(\bar{\nabla}_{X}\eta)(Y)=h(X,Y)-\frac{1}{2}A(X,Y,\eta^{\sharp}).$
Therefore, by Lemma 4.21, we conclude that
$\operatorname{div}_{h}(\eta^{\sharp})=-\star d\star\eta=-\varkappa(\eta^{\sharp})+n$
where $\star$ is the Hodge star operator on $(M,h)$. ∎
Corollary 4.23.
$\operatorname{Vol}_{h}(M)=\frac{1}{n}\int_{M}\varkappa\wedge\star\eta$
holds, where $\star$ is the Hodge star operator on $(M,h)$.
Proof.
Apply divergence theorem to the formula in Lemma 4.22. ∎
Corollary 4.24.
$\operatorname{Vol}_{h}(M)^{2}\leq\frac{1}{n^{2}}\left(\int_{M}|\varkappa|^{2}\operatorname{vol}_{h}\right)\left(\int_{M}|\eta|^{2}\operatorname{vol}_{h}\right)$
holds on $M$, with equality if and only if $\varkappa=k\cdot\eta$ for some positive real number $k\in\mathbb{R}$, that is, $h$ is a Cheng-Yau metric.
Proof.
Applying Cauchy-Schwarz inequality to the integral formula in Corollary 4.23, we immediately obatin the desired integral inequality. The equality holds exactly when there exists a real constant $k\in\mathbb{R}$ such that $\varkappa=k\cdot\eta$. By Corollary 4.23, we have
$\operatorname{Vol}_{h}(M)=\frac{1}{n}\int_{M}\varkappa\wedge\star\eta=\frac{k}{n}\int_{M}\eta\wedge\star\eta=\frac{k}{n}\int_{M}|\eta|^{2}\operatorname{vol}_{h}.$
Therefore the constant $k\in\mathbb{R}$ must be positive. ∎
A facile application of Corollary 4.24 is the following non-vanishing theorem for Riemann curvature tensors of compact Hessian manifolds of Koszul type.
Corollary 4.25.
Compact Hessian manifolds of Koszul type are not Riemannian flat.
Proof.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type, and let $\varkappa\in\Omega^{1}(M)$ be its first Koszul form.
Assume the contrary that $(M,h)$ is Riemannian flat. Then, by Theorem 2.44, we have that $\varkappa=0$. But Corollary 4.24 then yields $\operatorname{Vol}_{h}(M)\leq 0$, which is absurd. ∎
Remark 4.26.
The integral inequality stated in Corollary 4.24 is an analog of Heisenberg uncertainty principle in quantum mechanics: For a given dimension and volume, a $1$-form potential small in $L^{2}$-norm must be compensated by a first Koszul form large in $L^{2}$-norm, and vice-versa. This phenomenon indicates a highly non-trivial relation between a $1$-form potential and the first Koszul form of a Hessian manifold. We name this integral inequality as Koszul’s uncertainty principle, in honor of the great French mathematician Jean-Louis Koszul.
Remark 4.27.
On a compact Hessian manifold $((M,\nabla),h)$ of Koszul type, the integral inequality of Koszul’s uncertainty principle is sharp, and the equality is achieved exactly when $h$ is a Cheng-Yau metric. This reveals a deep connection with Theorem 4.14 and Proposition 4.19.
@@ 5 Topics in Low Dimensional Hessian Geometry
5.1 Exotic Potentials for the Euclidean Plane
Lemma 5.1.
Let $\Omega$ be a domain in $\mathbb{R}^{2}$, and let
Proof.
$F(x,y,z):=f(x/r,y/r,z/r)$
Proposition 5.2.
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
Proof.
Combine Corollary 4.4 with Lemma 5.1. ∎
Example 5.3.
$(2-d)H+x\frac{\partial H}{\partial x}+y\frac{\partial H}{\partial y}=0$
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a flat pseudo-Riemannian metric.
The following interesting example shows that nonlinear harmonic functions are Hessian potentials for the flat $1+1$ Minkowski plane $\mathbb{R}^{1,1}$.
Example 5.4.
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
We also suggest the following very concrete example of a Hessian potential for the flat Euclidean plane, which is constructed using the method introduced in Theorem 1.8.
Example 5.5.
$f(x,y):=\frac{x^{2}}{2y}+\frac{1}{4}\operatorname{log}(y)y.$
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}=\frac{1}{y}dx^{2}-\frac{2x}{y^{2}}dxdy+\frac{4x^{2}+y^{2}}{4y^{3}}dy^{2}$
$f_{xx}^{2}-4f_{xx}f_{yy}+4f_{xy}^{2}\equiv 0$
and then apply Proposition 5.2.
Before we close this section, we cite from [30] an example given by Dubrovin, to show that techniques from algebraic geometry, for example the Clebsch covariant of projective plane curves, might be useful for finding homogeneous potentials.
Example 5.6.
of $f$ is a flat metric on $\mathbb{R}^{3}$.
For more discussion on homogeneous polynomial potential of flat spaces, we refer to the work of B.Totaro [30].
@@ 5.2 Complete Hessian Surfaces
5.3 Compact Orientable Hessian 3-folds
@@
for every $z\in\mathbb{C}$, we obtain that $Q(\Phi(\gamma\cdot\varphi^{n}(\tau),\rho+n))=Q(\Phi(\tau,\rho))e^{2n}$, and hence
@@
In particular, we arrived at the following theorem:
Proposition 5.13.
Let $\Sigma_{g}$ be a closed orientable surface of genus $g>1$, and let $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ be a periodic mapping class. Then there exist a Riemannian metric $h$ on the mapping torus $M_{\varphi}$ of $\varphi$ and a flat torsion-free connection $D$ on the tangent bundle $T(M_{\varphi})\rightarrow M_{\varphi}$, such that $((M_{\varphi},D),h)$ is a closed and orientable radiant Hessian $3$-manifold of Koszul type with negative scalar curvature.
Under the scope of Thurston’s geometrization, the high genus examples in the proof of Proposition 5.13 are those locally modeled by $\mathbb{H}^{2}\times\mathbb{R}$. A natural next step is to construct similar examples for the genus one case. We will now prove that these examples indeed exist, and they all admit flat models.
Proposition 5.14.
Let $\varphi]\in\operatorname{Mod}(\mathbb{T}^{2})$ be a periodic mapping class. Then there exist a Riemannian metric $g$ on the mapping torus $M_{\varphi}$ of $\varphi$, such that $(M_{\varphi},g)$ is a closed orientable Bieberbach $3$-manifold.
Proof.
Recall from the arithmetic of $\operatorname{SL}_{2}(\mathbb{Z})\cong\operatorname{Mod}(\mathbb{T}^{2})$ that every finite order element of $\operatorname{SL}_{2}(\mathbb{Z})$ is conjugate to either $\pm\operatorname{Id}$ or one of the following matrices
$A_{1}=\left[\begin{array}[]{cc}0&-1\\ 1&-1\end{array}\right],\quad A_{2}=\left[\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right],\quad A_{3}=\left[\begin{array}[]{cc}0&-1\\ 1&1\end{array}\right]$
or their inverses. Now, it suffices to notice that $\pm\operatorname{Id},A_{2}$ are isometries of the rhombic elliptic curve $\mathbb{C}^{1}/(\mathbb{Z}\oplus i\mathbb{Z})$, and $A_{1},A_{3}$ are isometries of the hexagonal elliptic curve $\mathbb{C}^{1}/(\mathbb{Z}\oplus\omega\mathbb{Z})$, where $\omega\in\mathbb{C}$ is the $3$-rd primitive root of the unity. ∎
The examples constructed above are consistent with the results established earlier. As predicted by Theorem 2.46, for a closed orientable surface $\Sigma_{g}$ of genus $g\geq 1$, the mapping torus $M_{\varphi}$ of a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ is indeed a fiber bundle over $\mathbb{S}^{1}$. The fundamental group of $M_{\varphi}$ is a semidirect product of $\mathbb{Z}$ and $\pi_{1}(\Sigma_{g})$, which is indeed infinite and torsion free, as stated in Proposition 2.7 and Proposition 2.19.
However, given the nature of all examples that we constructed, one can be curious about whether there exists an example of closed orientable Hessian $3$-manifold that is a mapping torus of a mapping class of infinite order. We now prove that this is impossible, which leads to a classification theorem of closed Hessian $3$-manifolds.
Theorem 5.15.
Let $((M,\nabla),h)$ be a closed orientable Hessian $3$-manifold. Then there exists a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ such that $M$ is diffeomorphic to the mapping torus of $\varphi$, where $\Sigma_{g}$ is the closed orientable surface of genus $g\geq 1$.
Proof.
In the special case that $(M,h)$ is a flat Riemannian manifold, the desired theorem is proved by the classification of space groups.
From now on, we assume that $(M,h)$ is not flat. Then by Theorem 2.48, there exists a $\Sigma_{g}$-bundle $f\colon M\rightarrow\mathbb{S}^{1}$ for some closed orientable surface $\Sigma_{g}$ of genus $g$.
Since $M$ is aspherical by Corollary 2.14, we obtain that $\Sigma_{g}$ ditto. Therefore, in particular, $\Sigma_{g}$ is not homeomorphic to $\mathbb{S}^{2}$, and hence $g\geq 1$.
In particular, we have $\pi_{1}(M)=\pi_{1}(\Sigma_{g})\rtimes_{\varphi}\mathbb{Z}$ where $\varphi$ is the monodromy of the mapping torus $f\colon M\rightarrow\mathbb{S}^{1}$. Let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. Then there exists a complete Riemannian metric $g$ on $\mathbb{R}^{n}$ such that $(\tilde{M},\pi^{*}h)$ is isometric to the product of Riemannian manifolds $(\mathbb{R}^{n},g)$ and $(\mathbb{R},dt^{2})$. Denote by $\hat{\varphi}\colon\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ the lifting of $\varphi\colon\Sigma_{g}\rightarrow\Sigma_{g}$ to the universal covering spaces. Since $\pi_{1}(M)$ acts on $(\mathbb{R}^{n}\times\mathbb{R},g\oplus dt^{2})$ by isometry, in particular the smooth mapping
$\Phi\colon\mathbb{R}^{n}\times\mathbb{R}\rightarrow\mathbb{R}^{n}\times\mathbb{R},\quad(\textbf{x},t)\mapsto(\hat{\varphi}(\textbf{x}),t+1)$
is an isometry. Therefore we conclude that $\hat{\varphi}$ and hence $\varphi$ are isometries.
Denote $i\colon\Sigma_{g}\rightarrow M$ the inclusion map of a fiber of the fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$. Consider the subgroup $A$ of the isometry group $G$ of $(\Sigma_{g},i^{*}h)$ generated by $\varphi$, and denote by $\bar{A}$ its closure in $G$. Then it is readily seen that $\bar{A}$ is a compact abelian subgroup of $G$, and hence the identity component of $\bar{A}$ is a torus, Therefore, $\varphi$ is isotopic in $G$ to a group element of finite order.
Quod erat demonstrandum. ∎
@@
Combine Proposition 5.13 and Proposition 5.14, we obtain that the converse statement of Theorem 5.15 is also true: For every closed orientable surface $\Sigma_{g}$ of genus $g\geq 1$ and every periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$, there exist a Riemannian metric $h$ on the mapping torus $M_{\varphi}$ of $\varphi$ and a flat torsion-free connection $D$ on the tangent bundle $T(M_{\varphi})\rightarrow M_{\varphi}$, such that $((M_{\varphi},D),h)$ is a Hessian manifold. These thus gives a complete topological classification of closed orientable Hessian $3$-manifolds.
Theorem 5.15 is also consistent with earlier discoveries: Consider a closed orientable surface $\Sigma_{g}$ of genus $g$. Recall that pseudo-Anosov maps and Dehn twists are of infinite order. By Nielsen–Thurston classification, the mapping torus $M_{\varphi}$ of a pseudo-Anosov mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ admits a hyperbolic metric, which is forbidden by our previous Theorem 4.10. For the very same reason, we obtain that the nilpotent Heisenberg manifold cannot support any Hessian metric.
We close this section with a final remark: It is obvious that the product of two Hessian manifolds is naturally a Hessian manifold. However, the converse statement is far from being correct, as in the example above, the product of the circle $\mathbb{S}^{1}$ and a closed orientable surface $\Sigma_{g}$ of genus $g>1$ admits an affine structure, but $\Sigma_{g}$ itself does not admit any affine structure if $g\neq 1$ by the proof of Chern’s conjecture in dimension two. Therefore, in particular, the examples presented in this section are different to those constructed using Example 1.7 and Theorem 1.8.